OpenStudy (anonymous):

Write the Equation of the Line in standard form m= -4 (-2,-3)

OpenStudy (anonymous):

@Nnesha @triciaal

OpenStudy (sleepyjess):

Is that a coordinate?

Nnesha (nnesha):

okay so here is a equation y =mx +b standard form is Ax+By =C first replace m by -4 and replace x and y to solve for b same like last one

OpenStudy (anonymous):

yes

Nnesha (nnesha):

y = mx+b replace m by -4 then x and y by(x,y) and solve for b

OpenStudy (anonymous):

y = -4x+b

Nnesha (nnesha):

yeah right replace x and y by their values :

OpenStudy (anonymous):

-11 = b i think

Nnesha (nnesha):

perfect now you have y = -4x -11 last step to move -4x on left side so add 4x both side

OpenStudy (anonymous):

y + 4x = -11

Nnesha (nnesha):

yes right just write like this first x 4x +y =-11 yeah done :)

OpenStudy (anonymous):

ok thanks

Nnesha (nnesha):

\(\huge\color{pink}{{\rm good ~}~\rm job!!}\) my pleasure :)